# Sharp non-uniqueness for the Navier–Stokes equations in scaling critical spaces

### v1

@@ 1. Introduction
where $u=u(t,x):[0,\infty)\times\mathbb{R}^{n}\to\mathbb{R}^{n}$ is the unknown velocity field of the fluid, whereas $u_{0}=u_{0}(x):\mathbb{R}^{n}\to\mathbb{R}^{n}$ is a given initial datum satisfying $\operatorname{div}u_{0}=0$. We denote by $\mathbb{P}={\rm Id}+\nabla(-\Delta)^{-1}\operatorname{div}$ the Helmholtz projection.
In dimensions $n\geqslant 3$, Lions–Masmoudi [Lio-Mas] and Furioli–Rieusset–Terraneo [Fur-Rie-Ter-2000] proved that mild solutions to (1.1) are unique in the critical space $C([0,T);L^{n}(\mathbb{R}^{n}))$. The aim of this paper is to prove the sharpness of the choice of $L^{n}(\mathbb{R}^{n})$ in the sense that the uniqueness fails when one replaces $L^{n}(\mathbb{R}^{n})$ by a scaling critical space that is even slightly larger; for instance, relaxing $L^{n}(\mathbb{R}^{n})=\dot{F}_{n,2}^{0}(\mathbb{R}^{n})$ to a slightly larger Lizorkin–Triebel space $\dot{F}_{n,q}^{0}(\mathbb{R}^{n})$ with any $q>2$ leads to non-uniqueness of the mild solutions in $C([0,T);\dot{F}_{n,q}^{0}(\mathbb{R}^{n}))$. To achieve this, we show that the necessary and sufficient condition for the uniqueness of mild solutions to (1.1) in the scaling critical Besov spaces is given by $1\leqslant p<n$ with $1\leqslant q\leqslant\infty$, or $p=n$ with $1\leqslant q\leqslant 2$; see Figure 1 below. Moreover, for the non-unique case $p=n$ with $2<q\leqslant\infty$, or $n<p\leqslant\infty$ with $1\leqslant q\leqslant\infty$, we prove that the zero initial datum $u_{0}=0$ generates infinitely many small global solutions $u\in C([0,\infty);\dot{B}_{p,q}^{n/p-1}(\mathbb{R}^{n}))$, which asymptotically converge to non-trivial small steady states $U\in\dot{B}_{p,q}^{n/p-1}(\mathbb{R}^{n})$ solving the unforced stationary Navier–Stokes equations:
@@ 1.1. Known results related to our work
Before we state our main results precisely, we recall previous studies related to our work and mention the place of our work. By the Duhamel principle, we may formally rewrite (1.1) as the integral equation
$u(t)=e^{t\Delta}u_{0}-\int_{0}^{t}e^{(t-\tau)\Delta}\mathbb{P}\operatorname{div}(u(\tau)\otimes u(\tau))d\tau,$ (1.3)
@@
A solution of (1.3) is called a mild solution. It is natural to construct mild solutions in scaling critical spaces. Here, let us recall the notion of scaling criticality. If $u$ is a mild solution, then $u_{\lambda}(t,x)=\lambda u(\lambda^{2}t,\lambda x)$ also solves (1.1) for all $\lambda>0$. A Banach space $X$ is called a scaling critical space if ${\left\|u_{\lambda}(0,\cdot)\right\|}_{X}={\left\|u(0,\cdot)\right\|}_{X}$ for all $\lambda>0$. Typical examples of scaling critical spaces are as follows
$\dot{H}^{\frac{n}{2}-1}(\mathbb{R}^{n})\hookrightarrow L^{n}(\mathbb{R}^{n})\hookrightarrow L^{n,\infty}(\mathbb{R}^{n}),$ (1.4)
$\dot{B}_{p,q}^{\frac{n}{p}-1}(\mathbb{R}^{n})\hookrightarrow{\rm BMO}^{-1}(\mathbb{R}^{n})\qquad(1\leqslant p<\infty,\ 1\leqslant q\leqslant\infty).$ (1.5)
@@
This line of research originates from the pioneering work of Fujita–Kato [Fuj-Kato-64-ARMA], where they proved well-posedness in $\dot{H}^{n/2-1}(\mathbb{R}^{n})$, and was extended to $L^{n}(\mathbb{R}^{n})$ by Kato [Kato-84] and Giga–Miyakawa [Gig-Miy-85]. Moreover, Kozono–Yamazaki [Koz-Yam-94], Cannone–Planchon [Can-Pla-96], and Planchon [Pla-98] proved well-posedness in $\dot{B}_{p,q}^{n/p-1}(\mathbb{R}^{n})$ for $1\leqslant p<\infty$ and $1\leqslant q\leqslant\infty$. In the endpoint case $p=\infty$, Koch–Tataru [Koh-Tat-01] established well-posedness for small data in ${\rm BMO}^{-1}(\mathbb{R}^{n})$, however it was shown by [Bou-Pav-08, Yon-10, Wan-15] that (1.1) is ill-posed in $\dot{B}_{\infty,q}^{-1}(\mathbb{R}^{n})$ ($1\leqslant q\leqslant\infty$). We refer to [Kozono-Shimizu-2018, Takeuchi-2025-SIMA, Yam-2000, arXiv:2509.21272v2] for the forced Navier–Stokes flow in critical spaces.
@@
Next, we address the issue of uniqueness for the Navier–Stokes equations. This topic was brought to the forefront by Leray’s pioneering work [Leray-1934-Acta], which produced weak solutions for any $L^{2}$ initial datum but did not settle their uniqueness. It is well-known that the Leray–Hopf weak solutions are unique if they belong to the Ladyzhenskaya–Prodi–Serrin class $L^{p}(0,T;L^{q}(\mathbb{R}^{n}))$ with $n/q+2/p=1$; see [Serrin-1963, Masuda-1984, Escauriaza-Seregin-Sverak-2003, Kozono-Sohr-1996] for more detail. For the mild solutions, Lions–Masmoudi [Lio-Mas] and Furioli–Rieusset–Terraneo [Fur-Rie-Ter-2000] proved the uniqueness in the scaling critical class $C([0,T);L^{n}(\mathbb{R}^{n}))$ with $n\geqslant 3$, which corresponds to an endpoint $(p,q)=(\infty,n)$ of the Ladyzhenskaya–Prodi–Serrin class. Note that [Fur-Rie-Ter-2000] considered not only in $L^{n}(\mathbb{R}^{n})$ but also other critical spaces such as Morrey spaces. See also [Yam-2000] for the uniqueness of small mild solutions in $L^{\infty}(0,T;L^{n,\infty}(\mathbb{R}^{n}))$ and [Miu-2005] for the uniqueness in $C([0,T);bmo^{-1}(\mathbb{R}^{n}))\cap L^{\infty}_{\rm loc}((0,T);L^{\infty}(\mathbb{R}^{n}))$. The recent progress of this direction is given by Zhan [arXiv:2402.01174v2], where he considered the uniqueness of mild solutions in $C([0,T);\widetilde{L}^{n,\infty}(\mathbb{R}^{n}))$ with a new space $\widetilde{L}^{n,\infty}(\mathbb{R}^{n})$ which is larger than $L^{n}(\mathbb{R}^{n})$. In terms of the Besov spaces, the following proposition holds.
@@
The proof of Proposition 1.1 follows immediately by adopting the argument in [Fur-Rie-Ter-2000, Fuj-AIPHC, Iwabuchi-Okazaki-2025] in the case of (U1) and by the result of [Fur-Rie-Ter-2000] with the continuous embedding $\dot{B}_{n,q}^{0}(\mathbb{R}^{n})\hookrightarrow L^{n}(\mathbb{R}^{n})$ for the case of (U2). Note that Proposition 1.1 still holds if one replaces Besov spaces by Lizorkin–Triebel spaces as the continuous embedding $\dot{F}_{p,q}^{n/p-1}(\mathbb{R}^{n})\hookrightarrow L^{n}(\mathbb{R}^{n})$ for all $(p,q)$ satisfying (U1) or (U2).
@@
Finally, we briefly review works on non-uniqueness. One major direction originates from applying Nash’s convex integration to the 3D Euler equations on the torus (see, [DeLellis-Szekelyhidi-09, Isett-18, Daneri-Szekelyhidi-2017-ARMA, Lellis-Szekelyhidi-2013-Invent, Buckmaster-Lellis-Isett-Szekelyhidi-2015-AnnMath] for instance). Using an iterative scheme based on the Euler–Reynolds system, these works constructed Euler flows violating the energy conservation law and made substantial progress toward the Onsager conjecture. Although incorporating the viscous term was long thought to be fundamentally difficult, Buckmaster–Vicol [Buc-Vic-19] introduced the concept of intermittent convex integration and proved that uniqueness of weak solutions to the Navier–Stokes equations on $\mathbb{T}^{3}$ breaks down in the supercritical class $C([0,T);H^{\beta}(\mathbb{T}^{3}))$ for some $0<\beta\ll 1/2$. Cheskidov–Luo [Che-Luo-21] showed that an endpoint Ladyzhenskaya–Prodi–Serrin class $L^{2}(0,T;L^{\infty}(\mathbb{T}^{n}))$, is optimal with respect to the time integrability exponent: they constructed non-unique solutions in $L^{p}(0,T;L^{\infty}(\mathbb{T}^{3}))$ for $p<2$. They also considered another endpoint phenomenon in two dimensions and proved non-uniqueness in $C([0,T);L^{p}(\mathbb{T}^{2}))$ for $p<2$ (see [Che-Luo-22-AnnPDE]). Furthermore, Cheskidov–Zeng–Zhang [arXiv:2503.05692] proved that every initial datum in $H^{1/2}(\mathbb{T}^{3})$ generates infinitely many Navier–Stokes flows whose kinetic energy is continuous and monotonically decreasing in time. Remarkably, Coiculescu–Palasek [Coi-Pal-25] showed that even in ${\rm BMO}^{-1}(\mathbb{T}^{3})$, where well-posedness for small data is known, a certain large initial datum generates two smooth solutions. Their argument builds on ideas developed by Palasek [Pal-25-IMRN] for the dyadic Navier–Stokes model. Pushing this direction further, Cheskidov–Dai–Palasek [arXiv:2511.09556] observed that non-uniqueness in the form of instantaneous Type I blow-up may be generated from arbitrary $C^{\infty}$ solenoidal initial data. Another strategy to prove non-uniqueness, developed by Jia–Sverák [Jia-Sve-Invent, Jia-Sve-JFA], is to exploit the instability of certain self-similar solutions. A notable result in this direction is due to Albritton–Brué–Colombo [Alb-Bru-Col-AnnMath], who proved non-uniqueness of Leray–Hopf weak solutions to the forced Navier–Stokes equations. See Aoki–Maekawa [Aok-Mae] for the corresponding result on a two-dimensional half space. More recently, Hou–Wang–Yang [2509.25116v1] obtained an analogous statement in the unforced case $f\equiv 0$ by a computer-assisted proof.
@@ 1.2. Main results
The purpose of this paper is to show non-uniqueness of mild solutions to (1.1) in $C([0,T);\dot{B}_{p,q}^{n/p-1}(\mathbb{R}^{n}))$ for all $(p,q)$ satisfying neither (U1) nor (U2). Now, the precise statement of our main theorem reads as follows.
@@
Then, there exist constants $\varepsilon_{0}=\varepsilon_{0}(n,p,q)>0$ and $\eta=\eta(n,p,q)\geqslant 0$ such that for any $0<\varepsilon\leqslant\varepsilon_{0}$ and solenoidal initial datum $u_{0}\in\dot{B}_{p,q}^{n/p-1}(\mathbb{R}^{n})$ with ${\left\|u_{0}\right\|}_{\dot{B}_{p,q}^{n/p-1}}\leqslant\eta$, (1.1) possesses a global mild solution $u_{\varepsilon}\in C([0,\infty);\dot{B}_{p,q}^{n/p-1}(\mathbb{R}^{n}))$ satisfying
$\lim_{t\to\infty}{\left\|u_{\varepsilon}(t)-U_{\varepsilon}\right\|}_{\dot{B}_{p,q}^{\frac{n}{p}-1}}=0,$ (1.6)
where $U_{\varepsilon}\in\dot{B}_{p,q}^{n/p-1}(\mathbb{R}^{n})$ solves (1.2) and admits the estimate
$C^{-1}\varepsilon\leqslant{\left\|U_{\varepsilon}\right\|}_{\dot{B}_{p,q}^{\frac{n}{p}-1}}\leqslant C\varepsilon$ (1.7)
with some positive constant $C=C(n,p,q)$ independent of $u_{0}$ and $\varepsilon$.
@@
    whereas our non-unique solutions do not satisfy the auxiliary dissipative condition $\widetilde{L^{1}}(0,T;\dot{B}_{p,q}^{n/p+1}(\mathbb{R}^{n}))$.
@@
    The large-time behavior of global solutions to (1.1) with small initial data has been studied extensively; see [Pla-98, Nakasato-25-JDE, Gallagher-Iftimie-Planchon-2002, Gallagher-Iftimie-Planchon-2003, Takeuchi-2025, Watanabe-2025, Benameu-2015-JMAA] for instance. However, all of these results exploit the viscous dissipation and address only solutions that decay and converge asymptotically to zero. In contrast, the non-decaying global solutions $u_{\varepsilon}$ constructed in this paper are the first examples [^1] In the framework of weak solutions, the existence of non-dissipative solutions is widely known; see [Buc-Vic-19]. of non-dissipative unforced Navier–Stokes flows in critical spaces where the Stokes semigroup is strongly continuous. [^2] If this is not the case, solutions may not decay even at the linear level.
@@
    hold with some $\delta=\delta(p,q)\in(0,1)$. Thus, non-uniqueness for $u_{0}=0$ remains valid after replacing the Besov spaces by the corresponding Lizorkin–Triebel spaces. In particular, we see that enlarging $L^{n}(\mathbb{R}^{n})=\dot{F}_{n,2}^{0}(\mathbb{R}^{n})$ in [Lio-Mas, Fur-Rie-Ter-2000] even slightly to $\dot{F}_{n,q}^{0}(\mathbb{R}^{n})$ with $q>2$ instantly leads to non-uniqueness. Moreover, we obtain the breakdown of unconditional uniqueness for small solutions in $L^{\infty}(0,T;{\rm BMO}^{-1}(\mathbb{R}^{n}))$ since ${\rm BMO}^{-1}(\mathbb{R}^{n})=\dot{F}_{\infty,2}^{-1}(\mathbb{R}^{n})$. This is different from the phenomenon in [Coi-Pal-25], which concerns non-uniqueness of smooth solutions for large data in ${\rm BMO}^{-1}(\mathbb{T}^{3})$. See Figure 1.
@@
    Concerning non-uniqueness for steady flows, Luo [Luo-ARMA-2019] constructed nontrivial solutions belonging to $L^{2}(\mathbb{T}^{n})$ with $n\geqslant 4$. See [Rieusset-25-JFA] for two-dimensional case in $H^{-1}(\mathbb{T}^{2})$. Note that both spaces are supercritical and the three-dimensional case has been open. Our result appears to be the first to establish non-uniqueness of steady solutions in scaling critical spaces including three-dimensional case. Meanwhile, our method cannot be implemented in two dimensions. This obstruction essentially reduces to the well-known difficulty of the solvability for the two-dimensional stationary Navier–Stokes equations on $\mathbb{R}^{2}$.
@@
    As will be explained later, our construction of solutions relies on the structure provided by the classical Nash lemma, whereas we do not employ any of the periodic building blocks commonly used in convex-integration approaches to fluid equations, such as Beltrami or Mikado flows.

### v2

# arXiv:2602.19846v2[math.AP] 02 Mar 2026

@@ 1. Introduction
where $u=u(t,x):[0,\infty)\times\mathbb{R}^{n}\to\mathbb{R}^{n}$ is the unknown velocity field of the fluid, whereas $u_{0}=u_{0}(x):\mathbb{R}^{n}\to\mathbb{R}^{n}$ is a given initial datum satisfying $\operatorname{div}u_{0}=0$. We denote by $\mathbb{P}={\rm Id}+\nabla\operatorname{div}(-\Delta)^{-1}$ the Helmholtz projection onto divergence-free vector fields.
In dimensions $n\geqslant 3$, Lions–Masmoudi [Lio-Mas] and Furioli–Rieusset–Terraneo [Fur-Rie-Ter-2000] proved that mild solutions to (1.1) are unique in the critical space $C([0,T);L^{n}(\mathbb{R}^{n}))$. The aim of this paper is to prove the sharpness of the choice of $L^{n}(\mathbb{R}^{n})$ in the sense that the uniqueness fails as one replaces $L^{n}(\mathbb{R}^{n})$ by a scaling critical space that is even slightly larger; for instance, enlarging $L^{n}(\mathbb{R}^{n})=\dot{F}_{n,2}^{0}(\mathbb{R}^{n})$ to a slightly larger Lizorkin–Triebel space $\dot{F}_{n,q}^{0}(\mathbb{R}^{n})$ with any $q>2$ leads to non-uniqueness of the mild solutions in $C([0,T);\dot{F}_{n,q}^{0}(\mathbb{R}^{n}))$. To achieve this, we show that the necessary and sufficient condition for the uniqueness of mild solutions to (1.1) in the scaling critical Besov spaces is given by $1\leqslant p<n$ with $1\leqslant q\leqslant\infty$, or $p=n$ with $1\leqslant q\leqslant 2$; see Figure 1 below. Moreover, for the non-unique case $p=n$ with $2<q\leqslant\infty$, or $n<p\leqslant\infty$ with $1\leqslant q\leqslant\infty$, we prove that the zero initial datum $u_{0}=0$ generates infinitely many small global solutions $u\in C([0,\infty);\dot{B}_{p,q}^{n/p-1}(\mathbb{R}^{n}))$, which asymptotically converge to a non-trivial flow $U\in\dot{B}_{p,q}^{n/p-1}(\mathbb{R}^{n})$ solving the unforced stationary Navier–Stokes equations:
@@ 1.1. Known results related to our work
Before we state our main results precisely, we recall the previous studies related to our work. By the Duhamel principle, we may formally rewrite (1.1) as the integral equation
$u(t)=e^{t\Delta}u_{0}-\int_{0}^{t}e^{(t-\tau)\Delta}\mathbb{P}\operatorname{div}(u(\tau)\otimes u(\tau))d\tau.$ (1.3)
@@
A solution to (1.3) is called a mild solution of (1.1). In view of the Fujita–Kato principle, it is natural to construct mild solutions in scaling critical spaces. Here, let us recall the notion of scaling criticality. If $u$ is a mild solution, then $u_{\lambda}(t,x)=\lambda u(\lambda^{2}t,\lambda x)$ also solves (1.3) for all $\lambda>0$. A Banach space $X$ is called a scaling critical space if ${\left\|u_{\lambda}(0,\cdot)\right\|}_{X}={\left\|u(0,\cdot)\right\|}_{X}$ for all $\lambda>0$. Typical examples of scaling-critical spaces and their inclusion relations are as follows:
$\dot{H}^{\frac{n}{2}-1}(\mathbb{R}^{n})\hookrightarrow L^{n}(\mathbb{R}^{n})$ $\hookrightarrow L^{n,\infty}(\mathbb{R}^{n})$ (1.4)
$\hookrightarrow\dot{B}_{p,\infty}^{\frac{n}{p}-1}(\mathbb{R}^{n})\hookrightarrow{\rm BMO}^{-1}(\mathbb{R}^{n})\hookrightarrow\dot{B}_{\infty,\infty}^{-1}(\mathbb{R}^{n})$ (1.5)
with $n<p<\infty$ and
$\dot{B}_{p,q}^{\frac{n}{p}-1}(\mathbb{R}^{n})\hookrightarrow{\rm BMO}^{-1}(\mathbb{R}^{n})$ (1.6)
@@
with $1\leqslant p<\infty$ and $1\leqslant q\leqslant\infty$. This line of research originates from the pioneering work of Fujita–Kato [Fuj-Kato-64-ARMA], where they proved well-posedness in $\dot{H}^{n/2-1}(\mathbb{R}^{n})$, and was extended to $L^{n}(\mathbb{R}^{n})$ by Kato [Kato-84] and Giga–Miyakawa [Gig-Miy-85]. Moreover, Kozono–Yamazaki [Koz-Yam-94], Cannone–Planchon [Can-Pla-96], and Planchon [Pla-98] proved well-posedness in $\dot{B}_{p,q}^{n/p-1}(\mathbb{R}^{n})$ for $1\leqslant p<\infty$ and $1\leqslant q\leqslant\infty$. In the endpoint case $p=\infty$, Koch–Tataru [Koh-Tat-01] established well-posedness for small data in ${\rm BMO}^{-1}(\mathbb{R}^{n})$, however it was shown by [Bou-Pav-08, Yon-10, Wan-15] that (1.1) is ill-posed in $\dot{B}_{\infty,q}^{-1}(\mathbb{R}^{n})$ ($1\leqslant q\leqslant\infty$). See [IN] for more precise analysis on the well-posedness of (1.1) in critical Besov or Lizorkin–Triebel spaces. We refer to [Kozono-Shimizu-2018, Takeuchi-2025-SIMA, Yam-2000, arXiv:2509.21272v2] for the forced Navier–Stokes flow in critical spaces.
@@
Next, we address the issue of uniqueness for the Navier–Stokes equations. This topic was brought to the forefront by Leray’s pioneering work [Leray-1934-Acta], which produced weak solutions for any divergence-free initial datum in $L^{2}(\mathbb{R}^{3})$ but did not reveal their uniqueness. It is well-known that the Leray–Hopf weak solutions are unique if they belong to the Ladyzhenskaya–Prodi–Serrin class $L^{p}(0,T;L^{q}(\mathbb{R}^{n}))$ with $n/q+2/p=1$; see [Serrin-1963, Masuda-1984, Escauriaza-Seregin-Sverak-2003, Kozono-Sohr-1996] for more detail. For the mild solutions, Lions–Masmoudi [Lio-Mas] and Furioli–Rieusset–Terraneo [Fur-Rie-Ter-2000] proved uniqueness in the scaling critical class $C([0,T);L^{n}(\mathbb{R}^{n}))$ with $n\geqslant 3$, which corresponds to an endpoint $(p,q)=(\infty,n)$ of the Ladyzhenskaya–Prodi–Serrin class. See also [Yam-2000] for the uniqueness of small mild solutions in $L^{\infty}(0,T;L^{n,\infty}(\mathbb{R}^{n}))$ and [Miu-2005] for the uniqueness in $C([0,T);bmo^{-1}(\mathbb{R}^{n}))\cap L^{\infty}_{\rm loc}((0,T);L^{\infty}(\mathbb{R}^{n}))$. The recent progress of this direction is given by Zhan [arXiv:2402.01174v2], where he considered the uniqueness of mild solutions in $C([0,T);\widetilde{L}^{n,\infty}(\mathbb{R}^{n}))$ with a new space $\widetilde{L}^{n,\infty}(\mathbb{R}^{n})=\overline{L^{n,\infty}(\mathbb{R}^{n})\cap L^{\infty}(\mathbb{R}^{n})}^{L^{n,\infty}}$ which is strictly larger than the completion of $C_{c}^{\infty}(\mathbb{R}^{n})$ in $L^{n,\infty}(\mathbb{R}^{n})$. In terms of the scaling critical Besov spaces, the following proposition holds.
@@
The proof of Proposition 1.1 follows immediately by adopting the argument in [Fur-Rie-Ter-2000, Fuj-AIPHC, Iwabuchi-Okazaki-2025] in the case of (U1) and by the result of [Fur-Rie-Ter-2000] with the continuous embedding $\dot{B}_{n,q}^{0}(\mathbb{R}^{n})\hookrightarrow L^{n}(\mathbb{R}^{n})$ holds for the case of (U2). Note that Proposition 1.1 still holds if one replaces Besov spaces by Lizorkin–Triebel spaces as the continuous embedding $\dot{F}_{p,q}^{n/p-1}(\mathbb{R}^{n})\hookrightarrow L^{n}(\mathbb{R}^{n})$ for all $(p,q)$ satisfying (U1) or (U2).
@@
Finally, we briefly review works on non-uniqueness. One major direction originates from applying the Nash convex integration to the Euler equations on the torus (see, [DeLellis-Szekelyhidi-09, Isett-18, Daneri-Szekelyhidi-2017-ARMA, Lellis-Szekelyhidi-2013-Invent, Buckmaster-Lellis-Isett-Szekelyhidi-2015-AnnMath] for instance). Using an iterative scheme based on the Euler–Reynolds system via the geometric lemma, these works constructed Euler flows violating the energy conservation law and made substantial progress toward the Onsager conjecture. Although incorporating the viscous term was long thought to be fundamentally difficult, Buckmaster–Vicol [Buc-Vic-19] introduced the concept of intermittent convex integration and proved that uniqueness of weak solutions to the Navier–Stokes equations on $\mathbb{T}^{3}$ breaks down in the supercritical class $C([0,T);H^{\beta}(\mathbb{T}^{3}))$ for some $0<\beta\ll 1/2$. Cheskidov–Luo [Che-Luo-21] showed that an endpoint Ladyzhenskaya–Prodi–Serrin class $L^{2}(0,T;L^{\infty}(\mathbb{T}^{n}))$, is optimal with respect to the time integrability exponent; they constructed non-unique solutions in $L^{p}(0,T;L^{\infty}(\mathbb{T}^{3}))$ for $p<2$. In [Che-Luo-22-AnnPDE], they also considered another endpoint in two dimensions and proved non-uniqueness in $C([0,T);L^{p}(\mathbb{T}^{2}))$ for $p<2$. Furthermore, Cheskidov–Zeng–Zhang [arXiv:2503.05692] claimed that every initial datum in $H^{1/2}(\mathbb{T}^{3})$ generates infinitely many Navier–Stokes flows whose kinetic energy is continuous and monotonically decreasing in time. Remarkably, Coiculescu–Palasek [Coi-Pal-25] showed that even in ${\rm BMO}^{-1}(\mathbb{T}^{3})$, where well-posedness for small data is known, a certain large initial datum generates two smooth solutions. Their argument builds on ideas developed by Palasek [Pal-25-IMRN] for the dyadic Navier–Stokes model. Pushing this direction further, Cheskidov–Dai–Palasek [arXiv:2511.09556] observed that non-uniqueness, in the form of instantaneous Type I blow-up, may be generated from arbitrary $C^{\infty}$ solenoidal initial data. Another strategy to prove non-uniqueness is developed by Jia–Sverák [Jia-Sve-Invent, Jia-Sve-JFA], and is to exploit the instability of certain self-similar solutions. A notable result in this direction is due to Albritton–Brué–Colombo [Alb-Bru-Col-AnnMath], who proved non-uniqueness of Leray–Hopf weak solutions to the forced Navier–Stokes equations. See Aoki–Maekawa [Aok-Mae] for the corresponding result on a two-dimensional half space. More recently, Hou–Wang–Yang [2509.25116v1] announced an analogous statement in three-dimension with the unforced case $f\equiv 0$ by a computer-assisted proof.
@@ 1.2. Main results
The purpose of this paper is to show non-uniqueness of mild solutions to (1.1) in $C([0,T);\dot{B}_{p,q}^{n/p-1}(\mathbb{R}^{n}))$ for all $(p,q)$ satisfying neither (U1) nor (U2) in Proposition 1.1. Now, the precise statement of our main theorem reads as follows.
@@
Then, there exist a monotone decreasing sequence $\{\varepsilon_{m}\}_{m\in\mathbb{N}}\subset(0,1)$ with $\varepsilon_{m}\to 0$ and a constant $\eta=\eta(n,p,q)\geqslant 0$ such that for any $m\in\mathbb{N}$ and initial datum $u_{0}\in\dot{B}_{p,q}^{n/p-1}(\mathbb{R}^{n})$ with
${\left\|u_{0}\right\|}_{\dot{B}_{p,q}^{n/p-1}}\leqslant\eta,\qquad\operatorname{div}u_{0}=0,$ (1.7)
there exist a non-stationary global mild solution $u_{m}\in C([0,\infty);\dot{B}_{p,q}^{n/p-1}(\mathbb{R}^{n}))$ to (1.1) and a stationary solution $U_{m}\in\dot{B}_{p,q}^{n/p-1}(\mathbb{R}^{n})$ to (1.2) satisfying
$\lim_{t\to\infty}{\left\|u_{m}(t)-U_{m}\right\|}_{\dot{B}_{p,q}^{\frac{n}{p}-1}}=0,\qquad C^{-1}\varepsilon_{m}\leqslant{\left\|U_{m}\right\|}_{\dot{B}_{p,q}^{\frac{n}{p}-1}}\leqslant C\varepsilon_{m}$ (1.8)
with some constant $C=C(n,p,q)>0$ independent of $u_{0}$ and $m$.
@@
    whereas our non-unique solutions do not belong to the auxiliary dissipative space $\widetilde{L^{1}}(0,T;\dot{B}_{p,q}^{n/p+1}(\mathbb{R}^{n}))$. [^1] See Section 2 for the definition of Chemin–Lerner spaces $\widetilde{L^{\theta}}(0,T;\dot{B}_{p,q}^{s}(\mathbb{R}^{n}))$.
@@
    The large-time behavior of global solutions to (1.1) with small initial data in critical Besov spaces has been studied extensively; see [Pla-98, Nakasato-25-JDE, Gallagher-Iftimie-Planchon-2002, Gallagher-Iftimie-Planchon-2003, Takeuchi-2025, Watanabe-2025, Benameu-2015-JMAA] for instance. However, all of these results exploit the viscous dissipation and address only solutions that decay and converge asymptotically to zero. In contrast, the non-decaying global solutions $u_{\varepsilon}$ constructed in this paper are the first examples [^2] In the framework of weak solutions, the existence of non-dissipative solutions is widely known; see [Buc-Vic-19] for instance. of non-dissipative unforced Navier–Stokes flows in critical spaces in which the Stokes semigroup decays at temporal infinity.
@@
    hold with some $\delta=\delta(p,q)\in(0,1)$. Thus, non-uniqueness for $u_{0}=0$ remains valid if one replaces the Besov spaces by the corresponding Lizorkin–Triebel spaces. In particular, we see that enlarging $L^{n}(\mathbb{R}^{n})=\dot{F}_{n,2}^{0}(\mathbb{R}^{n})$ in [Lio-Mas, Fur-Rie-Ter-2000] even slightly to $\dot{F}_{n,q}^{0}(\mathbb{R}^{n})$ with $q>2$ leads to non-uniqueness. Moreover, we obtain the breakdown of unconditional uniqueness for small mild solutions in $L^{\infty}(0,T;{\rm BMO}^{-1}(\mathbb{R}^{n}))$ since ${\rm BMO}^{-1}(\mathbb{R}^{n})=\dot{F}_{\infty,2}^{-1}(\mathbb{R}^{n})$. This is different from the phenomenon in [Coi-Pal-25], which concerns non-uniqueness of smooth solutions for large data in ${\rm BMO}^{-1}(\mathbb{T}^{3})$. See Figure 1 below.
@@
    Concerning non-uniqueness for steady Navier–Stokes flows, Luo [Luo-ARMA-2019] constructed nontrivial solutions to (1.2) belonging to $L^{2}(\mathbb{T}^{n})$ with $n\geqslant 4$. See [Rieusset-25-JFA] for two-dimensional case in $H^{-1}(\mathbb{T}^{2})$. Note that both spaces are supercritical and the three-dimensional case has been open. Our result appears to be the first to establish non-uniqueness of steady solutions in scaling critical spaces. Moreover our result includes the three-dimensional case. Meanwhile, our method cannot be implemented in two dimensions. This obstruction essentially reduces to the well-known difficulty of the solvability for the two-dimensional stationary Navier–Stokes equations on $\mathbb{R}^{2}$.
@@
    As will be explained later, our construction of solutions relies on the structure provided by the classical Nash lemma, whereas we do not employ any of the periodic building blocks commonly used in convex integration approaches to fluid equations, such as Beltrami or Mikado flows.
  7. 7.
    Our proof works in essentially the same way if $\mathbb{R}^{n}$ is replaced by $\mathbb{T}^{n}$; see Remark 4.5.